# Fisher-Orthogonal Projection Methods for Natural Gradient Descent with Large Batches

### v2

@@ Introduction
The increasing scale of modern language models and vision transformers has made large mini-batch training a necessity. Modern GPUs offer extensive high-bandwidth memory (such as 192GB in AMD MI300X), and data centers often combine hundreds of these devices, enabling the efficient processing of tens of thousands of training examples per batch. This improves hardware utilization, reduces communication overhead, and accelerates training significantly. As batch sizes increase, the gradients become more deterministic. This reduces the stochastic noise that previously helped first-order optimizers such as SGD (Robbins and Monro 1951), Adam (Kingma and Ba 2014), and AdamW (Loshchilov and Hutter 2017) to explore flatter minima of the loss landscape. The loss of this noise  (Keskar et al. 2016; McCandlish et al. 2018) requires first-order methods to rely on smaller learning rates and stronger explicit regularization to maintain stability and generalization performance.
@@
Prior attempts to scale natural-gradient methods to large-batch training include SENG (Yang et al. 2020), which uses low-rank sketches but introducing new hyperparameters, and SP-NGD (Osawa et al. 2020), which relies on empirical-Fisher approximations and stale-statistic heuristics that require task-specific hyperparameter retuning. Adaptive batch-size schedules (Yao et al. 2018) improve throughput but still rely on heavily damped mean gradients. Although these methods reduce hardware requirements, their update rules remain fundamentally unchanged, still dominated by mean gradients and extensive hyperparameter tuning.
@@
Figure 1: 3D loss landscape of ResNet-18 with CIFAR-10 for batch size of 1024. Arrows represent the direction of the steps of different gradients. The green star is the smallest loss after updating the model based on different update directions.
@@
By construction, this ensures that $\langle g_{\text{avg}},g_{\text{diff}}^{\perp}\rangle_{F}=0$ (as described in Lemma 1 in the Appendix), meaning the new component contains only the information that is orthogonal in the Fisher geometry, which is already captured in the average gradient.
@@
Substituting this into the KL-norm expression we obtain a decomposition into three terms:
@@
where $Q=(F+\lambda I)^{-1}F(F+\lambda I)^{-1}$ acts as a curvature–weighted metric. In the large-damping case, where $\lambda\gg\Lambda_{i}$ for every Fisher eigenvalue $\Lambda_{i}$ that carries weight in $g$, the spectral factor $\frac{\Lambda_{i}}{(\Lambda_{i}+\lambda)^{2}}$ is proportional to $\Lambda_{i}/\lambda^{2}$. Therefore, the base term scales as $g^{\top}Qg\propto\lambda^{-2}.$ As shown in Appendix B, the two additional terms from FOP decay even more gently. Specifically:
@@
In this section, we rigorously evaluate FOP against both first‐order (SGD, AdamW) and second‐order (KFAC) baselines across four vision benchmarks: CIFAR-10 with ResNet-18, ImageNet-100 with T2T-ViT, ImageNet-1K with ResNet-50, and long-tailed CIFAR10-LT/100-LT with ResNet-32, demonstrating its fast convergence, large-batch scalability, and robustness under class imbalance. To ensure fair and rigorous comparisons, we evaluate all methods on several standard benchmarks: ResNet-18 on CIFAR-10, ResNet-50 on ImageNet-1k, and a Vision Transformer (ViT) on ImageNet-100. For each setting, we perform an extensive hyperparameter search for all optimizers. This includes tuning the learning rate, and for second-order methods like KFAC and FOP, also tuning the damping ratio $\lambda$. Following the linear scaling rule from (Goyal et al. 2017), the learning rate is scaled proportionally with the batch size. Additionally, the curvature update frequency for the second-order optimizer is reduced as the batch size increases, until it reaches a minimum threshold of 5 steps. To isolate the effect of our Fisher-orthogonal projection, FOP and KFAC share identical learning-rate schedules in every experiment. While our results highlight FOP’s advantages, we do not claim that FOP is a universally superior optimizer for all tasks and architectures. Rather, the evidence shows that augmenting natural gradient methods with a principled, geometry-aware variance component offers via FOP a robust and scalable path for second-order optimization in modern large-batch training scenarios. All experiments were performed on a single node with two AMD EPYC 9534 64-core CPUs and eight AMD MI300X GPUs. The implementations of KFAC and FOP rely on the ASDL package  (Osawa et al. 2023). Full details of the hyperparameter search space, such as learning rate, damping rate, and random seeding number, and a discussion about the memory overhead are provided in the Appendix.
@@
Batch Size Epochs / Time (s) and Speedup
(GPU) AdamW KFAC FOP (ours)
512 (1) 97 / 17 498.817\,498.8 42 / $8315.6$ 44 / $9535.7$
$\times 2.10$] $\times 1.84$]
1024 (2) 49 / $4603.6$ 41 / $4038.9$
$\times 3.80$] $\times 4.33$]
2048 (4) 57 / $2760.5$ 48 / $2535.4$
$\times 6.34$] $\times 6.90$]
4096 (8) 87 / $2715.3$ 49 / $1670.4$
$\times 6.45$] $\times 10.48$]
Table 2: Epoch and training time (in seconds) to reach 80.6% Top-1 accuracy for T2T-ViT on ImageNet-100. For KFAC and FOP, the bracketed numbers show the speedup factor relative to AdamW at the batch size of 512.
@@
Batch Size Epochs / Time (min) and Speedup
(GPU) SGD KFAC FOP (ours)
1024 (1) 71 / 2511.1 35 / 1336.5 32 / 1306.1
$\times 1.88$] $\times 1.92$]
2048 (2) – / – – / – 33 / 999.5
$\times 2.51$]
4096 (4) – / – – / – 34 / 514.9
$\times 4.88$]
8192 (8) – / – – / – 40 / 335.1
$\times 7.50$]
@@
To evaluate FOP at a much larger-scale dataset, we train ResNet-50 from scratch on ImageNet-1K (Deng et al. 2009) with running 3 different random seeds, to see if it can reach Top-1 test accuracy of 75.9%, which is a standard large-scale convolutional benchmark (Mattson et al. 2019). Following Yang et al. (2020); Liu et al. (2024), we run SGD for 80 epochs with a cosine learning-rate schedule, and we train both KFAC and FOP for 50 epochs using an exponential schedule on their learning rates.
@@ C.3. Setup of ImageNet-1K experiment
We implement ResNet50 following the official PyTorch example (Team 2017), based on the architecture proposed by He et al. (He et al. 2016). For all optimizers (SGD, Adam, KFAC, and FOP), we apply a linear warm-up strategy (Goyal et al. 2017) during the first 5 epochs. For the KFAC and FOP optimizers, the update intervals for the curvature matrix and its inverse are set to $\{800,1600,2400\}$ when using a batch size of $1024$. The total number of training epochs differs by optimizer: we train SGD and Adam for 80 epochs, while KFAC and FOP are trained for 50 epochs. SGD uses the cosine annealing learning rate schedule as described in Eq. C.2. Setup of ImageNet-100 experiment. In contrast, KFAC and FOP adopt an exponential decay learning rate schedule given by:

### v3

# Beyond the Mean: Fisher-Orthogonal Projection for Natural Gradient Descent in Large Batch Training

@@ Abstract
Codehttps://github.com/yishunlu-222/fop.git
Extended versionhttps://arxiv.org/abs/2508.13898
@@ Introduction
The increasing scale of modern language models and vision transformers has made large mini-batch training a necessity. Modern GPUs offer extensive high-bandwidth memory (such as 192GB in AMD MI300X), and data centers often combine hundreds of these devices, enabling the efficient processing of tens of thousands of training examples per batch. This improves hardware utilization, reduces communication overhead, and accelerates training significantly. As batch sizes increase, the gradients become more deterministic. This reduces the stochastic noise that previously helped first-order optimizers such as SGD (Robbins and Monro 1951), Adam (Kingma and Ba 2014), and AdamW (Loshchilov and Hutter 2017) to explore flatter minima of the loss landscape. The loss of this noise  (Keskar et al. 2016; McCandlish et al. 2018; You et al. 2019) requires first-order methods to rely on smaller learning rates and stronger explicit regularization to maintain stability and generalization performance.
@@
Prior attempts to scale natural-gradient methods to large-batch training include SENG (Yang et al. 2020), which uses low-rank sketches but introducing new hyperparameters, and SP-NGD (Osawa et al. 2020), which relies on empirical-Fisher approximations and stale-statistic heuristics that require task-specific hyperparameter retuning. Adaptive batch-size schedules (Yao et al. 2018; Shi et al. 2023) improve throughput but still rely on heavily damped mean gradients. Although these methods reduce hardware requirements, their update rules remain fundamentally unchanged, still dominated by mean gradients and extensive hyperparameter tuning.
@@
Figure 1: 3D loss landscape of training ResNet-18 with CIFAR-10 for batch size of 1024. This visualization relies on the method suggested in (Li et al. 2018). Arrows represent the direction of the steps of different gradients. The green star is the smallest loss after updating the model based on different update directions.
@@
By construction, this ensures that $\langle g_{\text{avg}},g_{\text{diff}}^{\perp}\rangle_{F}=0$ (as described in Lemma 1 in the Supplementary material in the extended version), meaning the new component contains only the information that is orthogonal in the Fisher geometry, which is already captured in the average gradient.
@@
Substituting this into the KL-norm expression, we obtain a decomposition into three terms:
@@
where $Q=(F+\lambda I)^{-1}F(F+\lambda I)^{-1}$ acts as a curvature–weighted metric. In the large-damping case, where $\lambda\gg\Lambda_{i}$ for every Fisher eigenvalue $\Lambda_{i}$ that carries weight in $g$, the spectral factor $\frac{\Lambda_{i}}{(\Lambda_{i}+\lambda)^{2}}$ is proportional to $\Lambda_{i}/\lambda^{2}$. Therefore, the base term scales as $g^{\top}Qg\propto\lambda^{-2}.$ As shown in Supplementary material in the extended version, the two additional terms from FOP decay even more gently. Specifically:
@@
In this section, we rigorously evaluate FOP against both first‐order (SGD, AdamW) and second‐order (KFAC) baselines across four vision benchmarks: CIFAR-10 with ResNet-18, ImageNet-100 with T2T-ViT, ImageNet-1K with ResNet-50, and long-tailed CIFAR10-LT/100-LT with ResNet-32, demonstrating its fast convergence, large-batch scalability, and robustness under class imbalance. To ensure fair and rigorous comparisons, we evaluate all methods on several standard benchmarks: ResNet-18 on CIFAR-10, ResNet-50 on ImageNet-1k, and a Vision Transformer (ViT) on ImageNet-100. For each setting, we perform an extensive hyperparameter search for all optimizers. This includes tuning the learning rate, and for second-order methods like KFAC and FOP, also tuning the damping ratio $\lambda$. Following the linear scaling rule from (Goyal et al. 2017), the learning rate is scaled proportionally with the batch size. Additionally, the curvature update frequency for the second-order optimizer is reduced as the batch size increases, until it reaches a minimum threshold of 5 steps. To isolate the effect of our Fisher-orthogonal projection, FOP and KFAC share identical learning-rate schedules in every experiment. While our results highlight FOP’s advantages, we do not claim that FOP is a universally superior optimizer for all tasks and architectures. Rather, the evidence shows that augmenting natural gradient methods with a principled, geometry-aware variance component offers via FOP a robust and scalable path for second-order optimization in modern large-batch training scenarios. All experiments were performed on a single node with two AMD EPYC 9534 64-core CPUs and eight AMD MI300X GPUs. The implementations of KFAC and FOP rely on the ASDL package  (Osawa et al. 2023). Full details of the hyperparameter search space, such as learning rate, damping rate, and random seeding number, and a discussion about the memory overhead are provided in the Supplementary material in the extended version.
@@
Optimizer Batch Size (GPU): Epochs / Time (min)
512 (1) 1024 (2) 2048 (4) 4096 (8)
AdamW 97 / $291.6$
KFAC 42 / $138.6$ 49 / $76.7$ 57 / $46.0$ 87 / $45.3$
$\times 2.10$] $\times 3.80$] $\times 6.34$] $\times 6.45$]
FOP (ours) 44 / $158.9$ 41 / $67.3$ 48 / $42.3$ 49 / $27.8$
$\times 1.84$] $\times 4.33$] $\times 6.90$] $\times 10.48$]
Table 2: Epoch and training time (in minutes) to reach 80.6% Top-1 accuracy for T2T-ViT on ImageNet-100. For KFAC and FOP, the bracketed numbers show the speedup factor relative to AdamW at the batch size of 512.
@@
Optimizer Batch Size (GPU): Epochs / Time (min)
1024 (1) 2048 (2) 4096 (4) 8192 (8)
SGD 71 / 2511.1 – / – – / – – / –
Shampoo 65 / 2523.1 – / – – / – – / –
LAMB 67 / 2492.7 – / – – / – – / –
KFAC 35 / 1336.5 – / – – / – – / –
[($\times 1.88$)]
FOP (ours) 32 / 1306.1 33 / 999.5 34 / 514.9 40 / 335.1
[($\times 1.92$)] [($\times 2.51$)] [($\times 4.88$)] [($\times 7.50$)]
@@
To evaluate FOP at a much larger-scale dataset, we train ResNet-50 from scratch on ImageNet-1K (Deng et al. 2009) with running 3 different random seeds, to see if it can reach Top-1 test accuracy of 75.9%, which is a standard large-scale convolutional benchmark (Mattson et al. 2019). Following Yang et al. (2020); Liu et al. (2024), we run SGD, Shampoo (Gupta, Koren, and Singer 2018), and LAMB (You et al. 2019) for 80 epochs with a cosine learning-rate schedule. We train both KFAC and FOP for 50 epochs using an exponential schedule on their learning rates, because in the SGD, Shampoo, and LAMB can’t reach the threshold of 75.9% within 50 epochs.
@@ Conclusion
Acknowledgments
This work was supported by UK Research and Innovation (UKRI) EP/T022205/1. The authors would also like to acknowledge support from the JADE 2.5 supercomputing service. The authors would also like to acknowledge the use of the University of Oxford Advanced Research Computing (ARC) facility (Richards 2015). The authors also acknowledge the use of resources provided by the Isambard-AI National AI Research Resource (AIRR). Isambard-AI is operated by the University of Bristol and is funded by the UK Government’s Department for Science, Innovation and Technology (DSIT) via UK Research and Innovation; and the Science and Technology Facilities Council [ST/AIRR/I-A-I/1023].
@@ C.3. Setup of ImageNet-1K experiment
We implement ResNet50 following the official PyTorch example (PyTorch Contributors 2025), based on the architecture proposed by He et al. (He et al. 2016). For all optimizers (SGD, Adam, KFAC, and FOP), we apply a linear warm-up strategy (Goyal et al. 2017) during the first 5 epochs. For the KFAC and FOP optimizers, the update intervals for the curvature matrix and its inverse are set to $\{800,1600,2400\}$ when using a batch size of $1024$. The total number of training epochs differs by optimizer: we train SGD and Adam for 80 epochs, while KFAC and FOP are trained for 50 epochs. SGD uses the cosine annealing learning rate schedule as described in Eq. C.2. Setup of ImageNet-100 experiment. In contrast, KFAC and FOP adopt an exponential decay learning rate schedule given by: